For the following market demand curve:


  1. Find marginal buyer’s WTP at Q = 15.
  2. Find CS for P = $35.
  3. Suppose P decreases to $10. How much will CS increase due to…

●  Buyers who enter the market

●  Existing buyers paying lower price.


1
Expert's answer
2021-04-27T07:05:37-0400

NOTE: Since the market demand curve is not provided, I will use the curve below to answer this question.





  1. The marginal buyer's WTP at Q=15 will be $25.
  2. CS at P= $35 will be calculated by finding the area of the triangle "(\frac{1}{2}\times b\times h)"

"= \frac{1}{2} \times 5\times 5"


"=\$12.5"

3 P decreases to 10,

  • Additional buyers will now fall between triangles A and B. Therefore, CS will be

"(\frac{1}{2} \times 10 \times 10)" "+ (\frac{1}{2} \times 15 \times 15)" "= \$162.5"

  • Increase in CS on the initial 10 units.

"10\times 15= \$150"